OpenStudy (anonymous):

two buses leave towns 924mi apart at the same time and travel toward each other. one bus travels 10mi/h slower than the other. if they meet in 6 hours, what is the rate of each bus?

Directrix (directrix):

Let x be speed of one bus Let x + 10 be speed of faster bus In 6 hours the first bus travels 6*x miles (speed times time = distance) In 6 hours the other bus travels 6*(x + 10) miles

Directrix (directrix):

6*x + 6* (x + 10) = 924 6*x + 6*x + 60 = 924 12*x = 864 x = 72 for speed of slower bus x + 10 = 82 for speed of faster bus

Directrix (directrix):

@Jenifa1771